In triangle ABC C=135 A=15 If AB=8 FIND AC AND BC USING 30 60 90 triangle

Grade:8

1 Answers

sonika p
6 years ago
A+B+C=180
15+135+B=180
B=30
using sine law
sinA/BC = sinB/AC = sinC/BA 
(sin135)/8 = sin15/BC
BC = (sin15)*8/(sin135) = 58.87
AC = (sin30)*8/(sin135) = 45.26